When Sara Whitehorse changed jobs, she rolled over the $7700.00 in her retirement account into two simple interest accounts. On one account, the annual simple interest rate is 8%, and on the second account, the annual simple interest is 6%. How much must be invested in each account if the two accounts earned the same amount of annual interest?

Answer by mananth(16946)

$x at 8%
7700-x at6%
6(7700-x)=8x
46200-6x=8x
14x=46200
/14
x=46200/14
x=3300
$3300 at 8%
Balance at6%